OpenStudy (anonymous):

Suppose that the funtions r and s are defined for all real numbers x as follows. r(x) = 3x^2 s(x) = 2x -- Write the expressions for (s-r)(x) and (s+r)(x) and evaluate (s*r)(2)

OpenStudy (anonymous):

(s-r)(x) = 2x-3x^2 , (s+r)(x)=2x+3x^2 , (s*r)(2) = 24

OpenStudy (anonymous):